A fly appears to be upside down and six times Its actual size when viewed through a converging lens with a focal length of 6.2 cm. How far is the fly from the lens? cm
Answer #1

Here ,

focal length , f = 6.2 cm

magnification , m = v/u

-6 = v/u ---(1)

Using lens formula

1/v -1/u = 1/6.2 ---(2)

solving the equations

v = 43.4 cm

u = - 7.23 cm

the distance of fly from the lens is 7.23 cm